two cells of emf E1=6 V and E2=5 v are joined in parallel with same polarity on same side, without any external load.if their internal resistances are r1=2 ohm and r2=3 ohm respectively, then terminal potential difference across any cell is

Grade:12

1 Answers

Vikas TU
9 years ago
Enet = 5.6 volt and I = 5.6/1.2 = 4.6 A I1 = 2/5 x 4.6 1.8 A terminal potential difference across any cell = 6 - 2I1 = 6- 2x2.7 = 6 - 5.4 = 0.6 volt plz approve.